OpenStudy (anonymous):

What is the solution of 10 equals the absolute value of 7 minus 3 x.?

OpenStudy (anonymous):

Ok so absolute value means you'll have two solution. 7-3x = 10 -3x = 3 ( move the 7 to the right hand side) x = -1 ( divide both sides by -3) and 7-3x = -10 -3x = -17 ( move 7 to the right hand side) x = -17/-3 = 17/3 = 5.6666... ( divide both sides by -3) so x = -1 and x = 5.6666 is a solution

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You're welcome :)